• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name moffit

    Log entry time 11:57:24 on October15,2002

    Entry number 88270

    This entry is a followup to: 88268


    keyword=Re-Setting Left arm momenta

    Ed Folts called and let us know that it's all right to put the magnets back to their momenta settings

    For left arm: set Q2, Q3 to 1600A (waited for 2 minutes at this point), then set p=2.01100. Dipole reached a stable current, but momentum isn't right

    *** Right arm will need to be cycled by S. Choi (page him an hour or so before ready to take LD2 data). Will take data with 14cm Dummy cell and left arm, until that time.

    Here's the resulting HAC screen:

    FIGURE 1